OpenStudy (anonymous):

find the domain of the rational function g(x)=x/x^3-216

OpenStudy (anonymous):

The denominator can't be zero ( undefined ) so that depends upon whether you mean\[g(x) = (x/x^{3}) - 216\]or\[g(x) = x/(x^3-216)\]hence the domain of the former is all real numbers except zero, the latter is all real numbers except 6.

OpenStudy (anonymous):

it was the second one